# Agent-Oriented Planning in Multi-Agent Systems

### v1

@@ A.1 Meta-agent Prompt
You are a planning agent. You are responsible for decomposing the given query into sub-tasks and choose the most suitable agent for each sub-task. Your main goal is to efficiently and accurately complete task planning based on the descriptions of agents provided, ensuring the coherence and quality of the sub-tasks.
Please output the sub-tasks and corresponding agents in the following json format: [{"task": task_description, "id": task_id, "name": name_of_agent, "reason": your_detailed_reason_for_the_choice, "dep":dependency_task_ids}]. In this format, "task" is the description of the sub-task, which will be used as the input of the chosen agents; "dep" denotes the id of the previous sub-task which generates a new resource relied by the current sub-task.
@@ A.2 Replan
The prompt for replanning is shown in Figure 4.
@@
You are a planning agent. You are preliminarily responsible for decomposing the given query into sub-tasks and choose the most suitable agent for each sub-task according to the following json format: [{"task": task_description, "id": task_id, "name": name_of_agent, "reason": your_detailed_reason_for_the_choice, "dep":dependency_task_ids}]. In this format, "task" is the description of the sub-task, which will be used as the input of the chosen agents; "dep" denotes the id of the previous sub-task which generates a new resource relied by the current sub-task.
@@
Figure 4: Prompt for replanning.
@@ A.3 Plan in Detail
You are a planning agent. You are preliminarily responsible for decomposing the given query into sub-tasks and choose the most suitable agent for each sub-task according to the following json format: [{"task": task_description, "id": task_id, "name": name_of_agent, "reason": your_detailed_reason_for_the_choice, "dep":dependency_task_ids}]. In this format, "task" is the description of the sub-task, which will be used as the input of the chosen agents; "dep" denotes the id of the previous sub-task which generates a new resource relied by the current sub-task.
@@ A.5 Detector
Appendix B Agents
The prompts for the code, math, search and commonsense agent are shown in Figure 8, 9, 10, 11 separately.
@@ Appendix C Constructing Training Dataset
The prompt for the scorer is shown in 12. The criteria for manual scoring are consistent with the standards outlined in the prompt provided to the scorer.
@@ 3. Completeness
We design a mapping function that converts different combinations of correctness, relevance, and completeness values into a final score, which is shown in Figure 13. Then, a score of 7 or higher is considered sufficiently high, while a score of 1 or lower is considered sufficiently low.
@@
Appendix D Evaluation
evaluate_prompt = ’’’
Question: %s
Ground-truth answer: %s
[Start of the prediction]
@@ Appendix D Evaluation
[End of the prediction]
Figure 14: Prompt for comparing the response with the ground truth.

### v2

@@ A.1 Meta-agent
You are a planning agent. You are responsible for decomposing the given query into sub-tasks and choosing the most suitable agent for each sub-task. Your main goal is to efficiently and accurately complete task planning based on the descriptions of agents provided, ensuring the coherence and quality of the sub-tasks.
Please output the sub-tasks and corresponding agents in the following JSON format: [{"task": task_description, "id": task_id, "name": name_of_agent, "reason": your_detailed_reason_for_the_choice, "dep":dependency_task_ids}]. In this format, "task" is the description of the sub-task, which will be used as the input of the chosen agent; "dep" denotes the id of the previous sub-task which generates a new resource relied by the current sub-task.
The available agents and the corresponding descriptions are: [code_agent: Generate code in Python for precise computations to solve the given task. math_agent: Answer math questions by reasoning step-by-step. search_agent: Call Bing Search API to obtain information regarding the given task. commonsense_agent: Answer the given question using commonsense reasoning.].
@@ A.2 Replan
The replan prompt is shown in Figure 4.
@@
You are a planning agent. You are preliminarily responsible for decomposing the given query into sub-tasks and choose the most suitable agent for each sub-task according to the following json format: [{"task": task_description, "id": task_id, "name": name_of_agent, "reason": your_detailed_reason_for_the_choice, "dep":dependency_task_ids}]. In this format, "task" is the description of the sub-task, which will be used as the input of the chosen agent; "dep" denotes the id of the previous sub-task which generates a new resource relied by the current sub-task.
@@
Figure 4: Replan prompt.
@@ A.3 Plan in Detail
You are a planning agent. You are preliminarily responsible for decomposing the given query into sub-tasks and choose the most suitable agent for each sub-task according to the following json format: [{"task": task_description, "id": task_id, "name": name_of_agent, "reason": your_detailed_reason_for_the_choice, "dep":dependency_task_ids}]. In this format, "task" is the description of the sub-task, which will be used as the input of the chosen agent; "dep" denotes the id of the previous sub-task which generates a new resource relied by the current sub-task.
@@ A.5 Detector
A.6 Evaluation
evaluate_prompt = ’’’
Question: %s
Ground-truth answer: %s
[Start of the prediction]
%s
[End of the prediction]
Figure 8: Prompt for comparing the response with the ground truth.
A.7 Agents
The prompts for the code, math, search and commonsense agent are shown in Figure 9, 10, 11, 12 separately. Specifically, the commonsense agent can generate commonsense information like the metal melting points and the atomic numbers of helium and hydrogen.
@@ A.7 Agents
Appendix B Fault Tolerance: Unexpected Changes in Agent Availability
We categorize changes in agent availability into two situations: those occurring during non-execution periods and those during execution.
  • During non-execution periods: AOP allows for the addition or removal of agents. The meta-agent can first broadcast a simple sync signal to determine agent availability. Only available agents are provided to the meta-agent for task allocation.
  • During execution: Changes in agent availability during execution indicate that an agent selected for a task may unexpectedly become unavailable, leading to the meta-agent not receiving a response to this sub-task. In such scenarios, the meta-agent is required to reassign the sub-task to another agent with similar capabilities or to further decompose the task (i.e., plan-in-detail).
@@ Appendix C Constructing Training Dataset
The prompt for the scorer is shown in Figure 13. The criteria for manual scoring are consistent with the standards outlined in the prompt provided to the scorer.
@@
We design a mapping function that converts different combinations of correctness, relevance, and completeness values into a final score, which is shown in Figure 14. Then, a score of 5 or higher is considered sufficiently high, while a score of 1 or lower is considered sufficiently low.
@@
Appendix D Experiments on More Datasets
D.1 Performance on DROP and IIRC
In addition to Husky-QA (Kim et al., 2024), we also conduct experiments on DROP (Dua et al., 2019) and IIRC (Ferguson et al., 2020). The experimental results are shown in Table 4, which indicate that AOP significantly outperforms the baselines, achieving at least a 5.0% improvement on both DROP and IIRC. These experimental results further confirm the effectiveness and advancements of AOP.
Table 4: Accuracy (%) comparisons between AOP and baselines on DROP and IIRC.
Method DROP IIRC
GPT-4o 23.0 33.0
CoT 26.0 35.0
Zero-Shot CoT 24.5 33.0
Meta-Agent 25.5 32.5
Meta-Agent: Traversal 27.5 34.0
REACT 29.0 36.0
HUSKY 28.0 36.5
AOP 34.0 41.5
AOP (reward model trained on Husky-QA) 32.0 39.0
D.2 Generalization Capability of the Reward Model
To evaluate the generalization capability of the reward model, we train a reward model based on Husky-QA and utilize it in experiments on the DROP and IIRC datasets. As shown at the bottom in Table 4, the reward model trained on one dataset demonstrates good generalization to other datasets (though it experiences a slight performance drop compared to the specifically trained reward model), achieving superior performance compared to baselines.
Appendix E Further Investigation of Expert Agents
E.1 Expert Agents
To investigate the effectiveness of expert agents powered by LLMs fine-tuned on task-specific datasets, we adopted Qwen2-Math-7B (Yang et al., 2024) as the backbone LLM for the math agent and DeepSeek-Coder-V2 (Zhu et al., 2024) as the backbone LLM for the code agent. The experimental results show that integrating these expert agents results in an additional 1.8% improvement on Husky-QA compared with using GPT-4. These results demonstrate the potential for further enhancements of AOP by incorporating more powerful expert agents.
E.2 Multiple Agents with Same Expertise
Based on AOP, we set up an experiment involving 4 different math agents, using GPT-3.5, GPT-4o, Qwen2-Math-7B, and Llama-3.2-3B, respectively. We instruct the multi-agent system to solve the complex math problem from MATH (Hendrycks et al., 2021). With AOP, the queries are decomposed into multiple queries and these agents run in parallel to resolve them. The experimental results are shown in Table 5, demonstrating the effectiveness (at least 6% improvements) of AOP.
Table 5: Accuracy (%) comparisons on MATH.
GPT-3.5 GPT-4o Qwen2-Math-7B-Instruct Llama-3.2-3B AOP
MATH 43 62 66 36 72
Appendix F Additional Experimental Results and Discussions
F.1 Quantitative Evaluation on Solvability, Completeness, and Non-redundancy
We conduct a quantitative evaluation to demonstrate the effectiveness of AOP in terms of solvability, completeness, and non-redundancy. Specifically, we compare the decomposed sub-tasks provided by AOP to those provided by GPT-4o. These sub-tasks are assessed by an LLM-based evaluator, providing binary scores for solvability, completeness, and non-redundancy, with scores of 1 for those that meet these principles and 0 for not. The experimental results (averaged scores), as summarized in Table 6, show that AOP achieves significant improvements.
Table 6: Quantitative Evaluation on Solvability, Completeness and Non-redundancy.
Solvability Completeness Non-redundancy
GPT-4o 0.763 0.822 0.986
AOP 0.938 0.969 0.993
F.2 Effect of Different Representation Approaches for Agent Description
In this study, we employ a combination of predefined natural language descriptions and representative works as the representations of agents’ descriptions. A natural language description can be manually provided or automatically generated, detailing the general and task-independent abilities of agents. While providing a comprehensive and detailed natural language description can be beneficial, it also requires effective prompt engineering. The representative (please refer to Sec. 4.3 for more details) works complement the natural language descriptions and are often task-dependent, allowing for continuous updates during execution.
We conducted experiments on Husky-QA to compare the effectiveness of different representation approaches. The results, shown in Table 7, indicate that using natural language descriptions achieves significantly superior performance compared to using only representative works, which motivates the majority of existing studies to adopt natural language descriptions. Besides, incorporating task-specific representative works on top of natural language descriptions leads to a further 11.9% performance boost, demonstrating the effectiveness of our proposed combined representation approach.
Table 7: Comparisons among Natural Language Descriptions and Representative Works.
Approaches Accuracy (%)
Natural Language Descriptions 31.8
Representative Works 16.8
Both 43.7
Besides, in this study, we employ the natural language descriptions adapted from Husky Kim et al. (2024) to ensure fair comparisons. To further investigate the effect of natural language descriptions, we provide two different versions: (i) LLM-generated: We instruct the LLM to generate the natural language descriptions based on how we describe the agents in Section 5.1; (ii) Expert-written: The natural language descriptions crafted by a human expert. The comparisons on Husky-QA are shown in Table 8. From the table, we can observe that using simple, fully automated natural language descriptions may somewhat affect the overall performance. Besides, incorporating more expert insights into the generation of natural language descriptions can provide an additional performance boost. These experimental results align with the current intuitive understanding of prompt engineering.
Table 8: Comparisons among Different Natural Language Descriptions.
Accuracy (%)
Descriptions adapted from Husky 43.7
LLM-generated descriptions 40.4
Expert-written descriptions 44.2
F.3 The Effectiveness of Detector in Handling Missing Dependencies
We conduct an experiment to evaluate the effectiveness of the detector in handling missing dependencies. To be more specific, the detector is instructed to determine whether there are any missing dependencies before a sub-task is assigned to an agent for execution. If such dependencies exist, the execution results of these dependent sub-tasks must also be provided as inputs. The adopted prompt is provided in Figure 15. If there exists any missing information, the meta-agent will add sub-tasks to complete the dependencies. The experiments conducted on Husky-QA show that, by enabling the detector to handle missing dependencies, AOP obtains a 2.5% performance increase in accuracy (from 43.7% to 46.2%).
dep_detect_prompt = ’’’
You are an intelligent detector tasked with determining whether the provided dependency information is sufficient to complete a given task. If the dependency information is insufficient, you will review all historical data to find supplemental information. If neither the dependency information nor the historical data is sufficient, you will identify and list the missing information. Besides, the information in the task or a given query can be viewed as available in the dependency information.
Input:
Task: {Task description}
Dependency Information: {Dependency information provided}
Historical Data: {Historical data (listed with serial numbers)}
Available query: {The query whose information can be viewd as available inthe dependency information.}
Requirements:
1. Assess Dependency Information: Evaluate if the provided dependency information is sufficient to complete the task. If sufficient, answer ***Yes***; if not, answer ***No***. Besides, if the answer is ***Yes***, there is no need to check the following requirements.
2. Review Historical Data: If the answer above is ***No***, check the historical data to identify any relevant entries that can supplement the task requirements. List the serial numbers of the relevant entries as ~~~numbers~~~.
3. Identify Missing Information: If the historical data also cannot supplement the missing details, explicitly list what information is still required to complete the task as \$\$\$required additional information\$\$\$.
Output Format (strictly follow this structure):
1. Is the dependency information sufficient: ***Yes***/***No***
2. Relevant information from historical data: ~~~numbers~~~ or ~~~None~~~
3. Missing information: \$\$\$Specific missing information\$\$\$ or \$\$\$None\$\$\$
Examples:
Input:
Task: Calculate the total population of China and the United States in 2022.
Dependency Information:
Chinas population in 2022 is 1.4 billion.
Historical Data:
1. Chinas population in 2022 is 1.4 billion.
2. The United States population in 2022 is 330 million.
3. Total world population in 2022: 8 billion.
Available query: If the populations of China and India were combined in 2022, how many countries with a population of 70,850,000 each could be formed from this total population without leaving anyone out?
Output:
1. Is the dependency information sufficient: ***No***
2. Relevant information from historical data: ~~~2~~~
3. Missing information: \$\$\$None\$\$\$
Input:
Task: Calculate the total population of China and the United States in 2022.
Dependency Information:
Chinas population in 2022 is 1.4 billion.
Historical Data:
1. Chinas population in 2022 is 1.4 billion.
2. Total world population in 2022: 8 billion.
Available query: If the populations of China and India were combined in 2022, how many countries with a population of 70,850,000 each could be formed from this total population without leaving anyone out?
Output:
1. Is the dependency information sufficient: ***No***
2. Relevant information from historical data: ~~~None~~~
3. Missing information: \$\$\$The United States population in 2022\$\$\$
Input:
Task: %s
Dependency Information:
@@ 3. Missing information: \$\$\$The United States population in 2022\$\$\$
Historical Data:
%s
Available query: %s
Output:
Figure 15: Prompt for the detector in handling missing dependencies.
F.4 The Limitations on the Number of Agents
AOP has no inherent limitations on the number of agents. However, as the meta-agent is powered by LLMs, increasing the number of agents is subject to the contextual window length restrictions of LLMs (descriptions of a large number of agents could exceed the LLM’s output length limit, such as 128K tokens), and the effectiveness of LLMs could be affected by the ability to handle long contexts.
In practical applications, it is important to consider that as the number of agents increases, there may be redundant and functionally similar agents, which motivates us to apply some extension strategies to handle. For example, agents can be grouped according to their abilities. When the meta-agent allocates sub-tasks, it can first select an agent group for each sub-task and then further choose the most suitable agent within the group. This strategy respects the LLM’s context window limits and enhances the accuracy of agent selection.